Number of atoms of oxygen in 24 g of o3?
1
Expert's answer
2020-05-14T14:33:34-0400

Solution:

Molar mass of O3 will be 3×Ar(O) = 3×16 = 48 g/mol.

Moles of O3 = Mass of O3 / Molar mass of O3

Moles of O3 = (24 g) / (48 g/mol) = 0.5 mol


According to Avogadro's number “Number of units in one mole of any substance is equal to 6.022×1023”. The units may be electrons, atoms, ions, or molecules.

Therefore,

Number of molecules of O3 = Avogadro's number × Moles of O3

Number of molecules of O3 = (6.022×1023) × (0.5 mol) = 3.011×1023 molecules

Number of molecules of O3 = 3.011×1023 molecules.

1 molecule of ozone contains 3 O atoms.

Hence, the number of O atoms is thrice the number of ozone molecules.

Number of atoms of oxygen = 3 × Number of molecules of O3

Number of atoms of oxygen = 3 × (3.011×1023) = 9.033×1023 atoms.


Answer: 9.033×1023 atoms of oxygen in 24 g of O3.